OpenStudy (anonymous):

heelp.!! two urns contain white balls and yellow balls. the first urn contains 2 white balls and 7 yellow balls and the second urn contains 10 white balls and 4 yellow balls. a ball is drawn at random from each urn. what is the probability that both balls are white?

OpenStudy (anonymous):

A)2/21 B)10/63 C)1/20 D)12/23

OpenStudy (anonymous):

put the number of white balls in the 1st urn over the total number of balls in the 1st urn 2/(2+7) = 2/9 put the number of white balls in the 2nd urn over the total number of balls in the 2nd urn 10/(10+4) = 10/14 or 5/7 multiply the 1st probability 2/9 times the 2nd probability 5/7

OpenStudy (anonymous):

ok i got 10/63??

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

yes---good job!

OpenStudy (anonymous):

thank you so much .!!!